OpenStudy (anonymous):

How do I do substitution method on the problem 3x-4y=10 2y=2x-6

OpenStudy (jdoe0001):

can you solve the 2nd one for "y"? what do you get?

OpenStudy (anonymous):

I know I have to get y by its self right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ok so ...let me try it

OpenStudy (anonymous):

um if i subtract -2y its going to cancel the y out on the equal sign

OpenStudy (jdoe0001):

yes, so you instead divide by 2 both sides

OpenStudy (jdoe0001):

\(\bf \cfrac{2y}{2}=\cfrac{2x-6}{2}\)

OpenStudy (anonymous):

whoa ok that was smart lol

OpenStudy (anonymous):

ok so 3x-4y=10 and y=2x-6

OpenStudy (anonymous):

now I do 3x-4(2x-6) = 10

OpenStudy (jdoe0001):

well... more like .. one sec

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} 3x-4{\color{red}{ y}}=10\\ 2y=2x-6\implies {\color{red}{ y}}=\cfrac{2x-6}{2}\implies\bf {\color{red}{ y}}=x-3 \end{array}\quad thus\\ \quad \\ 3x-4{\color{red}{ y}}=10\implies 3x-4({\color{red}{ x-3}})=10\)

OpenStudy (j2lie):

you have to substitute the 2y=2x-6 into the 3x. Like this 3(2y=2x-6)-4y=10.

OpenStudy (anonymous):

ohh i forgot the fraction part..thank you I see it nowww

OpenStudy (jdoe0001):

yw

OpenStudy (j2lie):

why do you have to divide 2x-6?